OpenStudy (anonymous):

Solve the equation by factoring. 8x2 + 22x = 21 x = (smaller value) x = (larger value)

OpenStudy (anonymous):

We have : \[8x^2+22x=21\iff 8x^2+22x-21=0 \\\iff (4x-3)(2x+7)=0 \\\iff 4x-3=0\quad or \quad 2x+7=0\] You can complete the rest as I think

OpenStudy (anonymous):

Thanks for your help. Any ideas what x equals?

OpenStudy (anonymous):

Show us your essays

OpenStudy (anonymous):

Essays? Not sure what you mean by that.

OpenStudy (anonymous):

Im sorry, I'm new to this website.

OpenStudy (anonymous):

I mean you can show us your work to find x

OpenStudy (anonymous):

I have no idea how to answer it. I'm trying to get through this algebra class by the skin of my teeth :(

OpenStudy (anonymous):

Thanks for your help

OpenStudy (anonymous):

I will help you ! We get two equations What is the 1st one ?